# Higher Lie characters and root enumeration in classical Weyl groups

### v1

@@ 1.1. Properness of root enumerators
Let $G$ be a finite group. For every integer $k$ define the $k$-th root enumerator $\varphi_{k}^{G}:G\to{\mathbb{Z}}$ by
$\varphi_{k}^{G}(g):=|\{h\in G\,:\,h^{k}=g\}|.$
Clearly, $\varphi_{k}$ is a class function on $G$. A classical result of Frobenius [6] (see also [13, Problem 4.7]) shows that, for any finite group $G$, all root enumerators are virtual characters, namely:
$\langle\varphi_{k}^{G},\chi\rangle\in{\mathbb{Z}}\qquad(\forall\chi\in{\rm Irr}(G),\,k\in{\mathbb{Z}}),$
@@
Which finite groups $G$ have the property that all $k$-th root enumerators $\varphi^{G}_{k}$ are proper $G$-characters? Equivalently, for which finite groups $G$ is $\langle\varphi_{k}^{G},\chi\rangle\geq 0$ for all $\chi\in{\rm Irr}(G)$ and $k\in{\mathbb{Z}}$?
Clearly, $\varphi^{G}_{0}$ and $\varphi^{G}_{1}$ are proper characters for any finite group $G$. A complete characterization of the finite groups for which $\varphi^{G}_{2}$ is a proper character was given in a celebrated work of Frobenius and Schur [7]; see also [13, §4]. In particular, if all the (complex) irreducible characters of $G$ can be obtained from real representations, then $\varphi^{G}_{2}$ is a multiplicity-free sum of all the irreducible characters.
This paper deals with the classical Weyl groups; for their classification into families ($A$, $B$ and $D$) see, e.g., [11, §2.4]. Kerber conjectured that for the classical Weyl group of type $A_{n-1}$, i.e., the symmetric group $S_{n}$ on $n$ letters, the $k$-th root enumerator is proper for all $n$ and $k$. This conjecture was proved by Scharf [18], presenting $\varphi_{k}^{S_{n}}$ as a multiplicity-free sum of higher Lie characters (defined below). For a different (non-constructive) proof, see [21, Ex. 7.69.c]. Scharf further proved [19, 2.4.18] that, for any finite group $G$, the $k$-th root enumerator $\varphi_{k}^{G\wr S_{n}}$ in the wreath product $G\wr S_{n}$ is a proper character, provided that $\varphi^{G}_{t}$ is a proper character for all the divisors $t$ of $k$. In particular, all $k$-th root enumerators in the classical Weyl group of type $B_{n}$, isomorphic to ${\mathbb{Z}}_{2}\wr S_{n}$, are proper. The problem remained open for the third family of classical Weyl groups, those of type $D_{n}$ (for general $k$).
@@
In fact, we obtain an explicit expression of each root enumerator in $D_{n}$ as half the sum of restrictions (from $B_{n}$) of appropriate higher Lie characters of type $B$. On the way, we obtain an explicit expression of each root enumerator in $B_{n}$ as a sum of higher Lie characters. This is a new proof of the properness of root enumerators in $B_{n}$.
Combining the above mentioned results of Scharf with Theorem 1.2 we conclude the following.
@@ 1.2. Higher Lie characters and their role
(Higher Lie characters) Let $G$ be a finite group. A family of higher Lie characters (HLC) for $G$ is a family $\{\psi^{C}\,:\,C\in\operatorname{Conj}(G)\}$ of characters of $G$, indexed by conjugacy classes, satisfying the following conditions.
@@
    For each $C\in\operatorname{Conj}(G)$ there is an element $g\in C$ such that $\psi^{C}$ is induced from a linear character of the centralizer $Z_{G}(g)$.
@@
Such characters were constructed for conjugacy classes of involutions $(k=2)$ in $B_{n}$ by Baddeley [2] and by Scharf [19, §3.1]. The construction was generalized by Scharf [19, 2.4.21 Folgerung] to conjugacy classes of prime order (i.e., for $k$ prime) in wreath products. For general conjugacy classes in a wreath product, Scharf defined characters induced from linear characters on subgroups which are not necessarily centralizers.
In Section 4 we construct higher Lie characters $\psi^{C}$ for all conjugacy classes $C$ in $B_{n}\cong{\mathbb{Z}}_{2}\wr S_{n}$; see Definition 4.2 below. We prove the following.
@@
For any integer $k$ and positive integer $n$,
$\varphi_{k}^{B_{n}}=\sum_{C\in\operatorname{Conj}(B_{n})\,:\,C^{k}=\{1\}}\psi^{C}.$
@@
$\sum_{C\in\operatorname{Conj}(B_{n})\,:\,C^{k}=\{1\}}\psi^{C}(y)+\sum_{C\in\operatorname{Conj}(B_{n})\,:\,C^{k}=\{w_{0}\}}\psi^{C}(y)=\begin{cases}2\varphi_{k}^{D_{n}}(y),&\text{if }y\in D_{n};\\ 0,&\text{otherwise,}\end{cases}$
@@
If $n$ is odd, we can define higher Lie characters $\psi_{D_{n}}^{C}$ for all conjugacy classes $C$ in $D_{n}$; see Definition 6.10. If $n$ is even, there is no family of induced characters that satisfies Definition 1.4. However, Definition 6.10 can still be used for odd $k$. We then have the following improvement of Theorem 1.6.
@@
Let $k$ and $n\geq 2$ be integers such that $n$ or $k$ is odd. Then
$\varphi_{k}^{D_{n}}=\sum_{C\in\operatorname{Conj}(D_{n})\,:\,C^{k}=\{1\}}\psi_{D_{n}}^{C}.$
A Gelfand model of a finite group $G$ is a multiplicity-free sum of its irreducible characters. By the remarks above, if all the irreducible characters of $G$ come from real representations, then a Gelfand model for $G$ yields the character $\varphi_{2}^{G}$. The problem of constructing Gelfand models for groups was introduced by Bernstein, Gelfand and Gelfand [4]. A construction of a Gelfand model for the symmetric group $S_{n}$, as a multiplicity-free sum of characters induced from centralizers of involutions, was presented by Klyachko [14] and by Inglis, Richardson and Saxl [12]. Such a construction is called an involution model. Baddeley [2] generalized this construction to Weyl groups of types $B_{n}$ (for all $n$) and $D_{n}$ (for odd $n$), and noted [3, Proposition 4.8.1] the non-existence of such a construction for Weyl groups of type $D_{n}$ when $n$ is even. Theorem 1.6 can be used to construct a double Gelfand model for $D_{n}$, namely, a character of $D_{n}$ in which every irreducible character appears exactly twice, see Theorem 7.3 below.
The rest of the paper is organized as follows. Section 2 contains necessary background and notation. In Section 3 we compute a generating function (Theorem 3.1) for root enumerators in $B_{n}$. Section 4 is devoted to higher Lie characters of type $B$, including an explicit definition (Definition 4.2) and a proof of a generating function (Theorem 4.3) for their values. In Section 5 we compare the two generating functions and prove Theorem 1.5, restated as Theorem 5.3. In Section 6 we consider the groups $D_{n}$ and prove Theorems 1.6, 1.2 and 1.7. The paper concludes with applications and open problems regarding Gelfand models (Subsection 7.1), index $2$ subgroups (Subsection 7.2), and related quasisymmetric functions (Subsection 7.3).
Acknowledgements
The authors thank Axel Hultman, Eric Marberg and Richard Stanley for useful discussions and references.
@@ 2.2. The Weyl groups of types $A$, $B$ and $D$
In this paper we use an explicit description of the conjugacy classes in $B_{n}$ (but do not need the corrsponding description in $D_{n}$). A bipartition ${\underline{\lambda}}=(\lambda^{+},\lambda^{-})$ of a positive integer $n$, denoted ${\underline{\lambda}}\vdash n$, is a pair of partitions of total size $n$. Conjugacy classes in $B_{n}$ are in one-to-one correspondence with bipartitions of $n$. Each element $w\in B_{n}$, viewed as a permutation of the set [±$[\pm n]$, can be written as a product of disjoint cycles of total length $2n$. Moreover, if $c=(a_{1},a_{2},\ldots,a_{\ell})$ is such a cycle then so is $\bar{c}=(-{a}_{1},-{a}_{2},\ldots,-{a}_{\ell})$, and either $c$ and $\bar{c}$ are disjoint or else $c=\bar{c}$. In the former case, the product $c\bar{c}$ is said to be a positive cycle of $w$, of length $\ell$; otherwise $\ell$ is necessarily even and $c=\bar{c}$ is said to be a negative cycle of $w$, of length $\ell/2$. Then the pair of partitions $(\lambda^{+},\lambda^{-})$ for which the parts of $\lambda^{+}$ (respectively, $\lambda^{-}$) are the lengths of the positive (respectively, negative) cycles of $w$ is a bipartition of $n$, called the signed cycle type of $w$. Two elements of $B_{n}$ are conjugate if and only if they have the same signed cycle type. We denote by $C_{\underline{\lambda}}$ the conjugacy class of elements of $B_{n}$ of signed cycle type ${\underline{\lambda}}=(\lambda^{+},\lambda^{-})$.
@@ 2.3. Higher Lie characters of type $A$
$\psi^{\lambda}:=\omega^{\lambda}\uparrow_{Z_{\lambda}}^{S_{n}}.$
@@
$\varphi^{S_{n}}_{k}=\sum_{\lambda\vdash_{k}n}\psi^{\lambda}.$
@@ 3. Generating function for root enumerators of type $B$
In this section we compute, for any integer $k$, an explicit generating function for the values of the $k$-th root enumerator $\varphi_{k}^{B_{n}}$ on arbitrary elements of $B_{n}$. The result is expressed in terms of the $h$-th root enumerators $\varphi_{h}^{{\mathbb{Z}}_{2}}$, corresponding to the cyclic group ${\mathbb{Z}}_{2}$ of order 2 (written multiplicatively as $\{+1,-1\}$) for various divisors $h$ of $k$.
@@ 3.1. Main result
$\sum_{n\geq 0}\sum_{{\underline{\mu}}\vdash n}\frac{\varphi_{k}^{B_{n}}({\underline{\mu}})}{|Z_{\underline{\mu}}|}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}=\exp\left(\sum_{j\geq 1}\sum_{\theta\in{\mathbb{Z}}_{2}}\sum_{\begin{subarray}{c}h|k\\ \gcd(h,j)=1\end{subarray}}\frac{t_{j,\theta}^{jk/h}}{2jk/h}\cdot\varphi_{h}^{{\mathbb{Z}}_{2}}(\theta)\right).$
@@
$\sum_{n\geq 0}\sum_{\mu\vdash n}\frac{\varphi_{k}^{S_{n}}(\mu)}{|Z_{\mu}|}{\underline{t}}^{{\underline{c}}(\mu)}=\exp\left(\sum_{j\geq 1}\sum_{\begin{subarray}{c}h|k\\ \gcd(h,j)=1\end{subarray}}\frac{t_{j}^{jk/h}}{jk/h}\right).$
@@
In order to compute the size of the conjugacy class of $y$, first write an arbitrary permutation of the numbers $1,\ldots,jb$. Interpret the first $j$ numbers as forming a cycle, and assign to each of the numbers an element of ${\mathbb{Z}}_{2}$ such that the ${\mathbb{Z}}_{2}$-class of the whole cycle is $\theta$. This can be done in $|{\mathbb{Z}}_{2}|^{j-1}$ ways. Treat similarly the other cycles, each having length $j$ and ${\mathbb{Z}}_{2}$-class $\theta$. Noting that any permutation of the $b$ cycles, as well as an arbitrary cyclic shift of each cycle, yield the same element of $B_{jb}$, it follows that the size of the conjugacy class of $y$ in $B_{jb}$ is
@@
so that $|R_{h,\theta}|=\varphi^{{\mathbb{Z}}_{2}}_{h}(\theta)$.
@@
$\sum_{{\underline{\lambda}}\vdash n}\sum_{\begin{subarray}{c}x\in C_{\underline{\lambda}}\\ x^{k}=y\end{subarray}}1=\varphi_{k}^{B_{n}}(y).$
Recall that $|R_{h,\theta}|=\varphi_{h}^{{\mathbb{Z}}_{2}}(\theta)$ and $|Z_{\underline{\mu}}|=|B_{n}|/|C_{\underline{\mu}}|$. Denoting ${\underline{t}}^{{\underline{c}}({\underline{\mu}})}:=\prod_{j,\theta}t_{j,\theta}^{jb_{j,\theta}}$, Theorem 3.6 reduces to Theorem 3.1, presenting the (unrefined) generating function for the $k$-th root enumerator in $B_{n}$. ∎

### v2

@@ 1.1. Properness of root enumerators
Let $G$ be a finite group. For every integer $k$ define the $k$-th root enumerator $\rho_{k}^{G}:G\to{\mathbb{Z}}$ by
$\rho_{k}^{G}(g):=|\{h\in G\,:\,h^{k}=g\}|.$
Clearly, $\rho_{k}^{G}$ is a class function on $G$. A classical result of Frobenius [6] (see also [13, Problem 4.7]) shows that, for any finite group $G$, all root enumerators are virtual characters, namely:
$\langle\rho_{k}^{G},\chi\rangle\in{\mathbb{Z}}\qquad(\forall\chi\in{\rm Irr}(G),\,k\in{\mathbb{Z}}),$
@@
Which finite groups $G$ have the property that all $k$-th root enumerators $\rho^{G}_{k}$ are proper $G$-characters? Equivalently, for which finite groups $G$ is $\langle\rho_{k}^{G},\chi\rangle\geq 0$ for all $\chi\in{\rm Irr}(G)$ and $k\in{\mathbb{Z}}$?
Clearly, $\rho^{G}_{0}$ (the regular character) and $\rho^{G}_{1}$ (the trivial character) are proper characters for any finite group $G$. A complete characterization of the finite groups for which $\rho^{G}_{2}$ is a proper character was given in a celebrated work of Frobenius and Schur [7]; see also [13, §4]. In particular, if all the (complex) irreducible characters of $G$ can be obtained from real representations, then $\rho^{G}_{2}$ is a multiplicity-free sum of all the irreducible characters.
This paper deals with the classical Weyl groups; for their classification into families ($A$, $B$ and $D$) see, e.g., [11, §2.4]. Kerber conjectured that, for the classical Weyl group of type $A_{n-1}$ (i.e., the symmetric group $S_{n}$ on $n$ letters), the $k$-th root enumerator is proper for all $n$ and $k$. This conjecture was proved by Scharf [18], presenting $\rho_{k}^{S_{n}}$ as a multiplicity-free sum of higher Lie characters (to be defined below). For a different (non-constructive) proof, see [21, Ex. 7.69.c]. Scharf further proved [19, 2.4.18] that, for any finite group $G$, the $k$-th root enumerator $\rho_{k}^{G\wr S_{n}}$ in the wreath product $G\wr S_{n}$ is a proper character, provided that $\rho^{G}_{t}$ is a proper character for all the divisors $t$ of $k$. In particular, all $k$-th root enumerators in the classical Weyl group of type $B_{n}$, isomorphic to ${\mathbb{Z}}_{2}\wr S_{n}$, are proper. The problem remained open, for general $k$, for the third family of classical Weyl groups, those of type $D_{n}$.
@@
Moreover, the root enumerators can be expressed as sums of higher Lie characters: Each root enumerator in $D_{n}$ is explicitly expressed as half the sum of restrictions (from $B_{n}$) of appropriate higher Lie characters of type $B$ (Theorem 1.6). We also obtain an explicit expression of each root enumerator in $B_{n}$ as a sum of higher Lie characters of type $B$ (Theorem 1.5). This gives a new proof of Scharf’s result that root enumerators in $B_{n}$ are proper.
Combining Scharf’s results with Theorem 1.2, we conclude the following.
@@ 1.2. Higher Lie characters and their role
(Higher Lie characters) Let $G$ be a finite group. A family of higher Lie characters (HLC) for $G$ is a family $\{\psi_{G}^{C}\,:\,C\in\operatorname{Conj}(G)\}$ of characters of $G$, indexed by conjugacy classes, satisfying the following conditions.
@@
    For each $C\in\operatorname{Conj}(G)$ there is an element $g\in C$ such that $\psi_{G}^{C}$ is induced from a linear character of the centralizer $Z_{G}(g)$.
@@
Such characters were constructed for conjugacy classes of involutions $(k=2)$ in $B_{n}$ by Baddeley [2] and by Scharf [19, §3.1]. The construction was generalized by Scharf [19, 2.4.21 Folgerung] to conjugacy classes of prime order in wreath products. For general conjugacy classes in a wreath product, Scharf defined characters induced from linear characters on subgroups which are not necessarily centralizers.
In Section 4 we construct higher Lie characters of type $B$. These characters (see Definition 4.2 below) are induced from linear characters of centralizers, and therefore satisfy Definition 1.4(a). By the following theorem, they also satisfy Definition 1.4(b).
@@
For any integer $k$ and positive integer $n$, the family $\{\psi_{B_{n}}^{C}\,:\,C\in\operatorname{Conj}(B_{n})\}$ of $B_{n}$-characters from Definition 4.2 satisfies
$\rho_{k}^{B_{n}}=\sum_{C\in\operatorname{Conj}(B_{n})\,:\,C^{k}=\{1\}}\psi_{B_{n}}^{C}.$
@@
$\sum_{C\in\operatorname{Conj}(B_{n})\,:\,C^{k}=\{1\}}\psi_{B_{n}}^{C}(y)+\sum_{C\in\operatorname{Conj}(B_{n})\,:\,C^{k}=\{w_{0}\}}\psi_{B_{n}}^{C}(y)=\begin{cases}2\rho_{k}^{D_{n}}(y),&\text{if }y\in D_{n};\\ 0,&\text{otherwise,}\end{cases}$
@@
Remark 1.7.
By Theorem 1.5, the first sum on the LHS of the formula in Theorem 1.6 is equal to $\rho_{k}^{B_{n}}(y)$. In other words, $\rho_{k}^{D_{n}}(y)$ is half the sum of $\rho_{k}^{B_{n}}(y)$ and a correcting term. When both $k$ and $n$ are even, there does not seem to be a direct connection between the two root enumerators. However, in the following two cases there is a simple connection:
  1. 1.
    For $k$ even and $n$ odd, the second sum is zero. It follows that, in this case, $\rho_{k}^{D_{n}}(y)=\rho_{k}^{B_{n}}(y)/2$ for any $y\in D_{n}$. Indeed, in this case, for $x\in B_{n}$: $x^{k}=y$ if and only if $(xw_{0})^{k}=y$, and |$|\{x,xw_{0}\}\cap D_{n}|=1$.
  2. 2.
    For odd $k$ and any $n$, a parity argument shows that $\rho_{k}^{D_{n}}$ is the restriction (to $D_{n}$) of $\rho_{k}^{B_{n}}$. Indeed, one can verify that, in this case, the two sums in Theorem 1.6 are equal.
Furthermore, if $n$ is odd, we can define higher Lie characters $\psi_{D_{n}}^{C}$ for all conjugacy classes $C$ in $D_{n}$; see Definition 6.10. If $n$ is even then there is no family of induced characters that satisfies Definition 1.4, but Definition 6.10 can still be used for odd $k$. We therefore have the following improvement of Theorem 1.6.
@@
Let $k$ and $n\geq 1$ be integers which are not both even. Then
$\rho_{k}^{D_{n}}=\sum_{C\in\operatorname{Conj}(D_{n})\,:\,C^{k}=\{1\}}\psi_{D_{n}}^{C}.$
See Subsection 6.5 below.
A Gelfand model of a finite group $G$ is a multiplicity-free sum of its irreducible characters. By the remarks above, if all the irreducible characters of $G$ come from real representations, then a Gelfand model for $G$ yields the character $\rho_{2}^{G}$. The problem of constructing Gelfand models for groups was introduced by Bernstein, Gelfand and Gelfand [4]. A construction of a Gelfand model for the symmetric group $S_{n}$, as a multiplicity-free sum of characters induced from centralizers of involutions, was presented by Klyachko [14] and by Inglis, Richardson and Saxl [12]. Such a construction is called an involution model. Baddeley [2] generalized this construction to Weyl groups of types $B_{n}$ (for all $n$) and $D_{n}$ (for odd $n$), and noted [3, Proposition 4.8.1] the non-existence of such a construction for Weyl groups of type $D_{n}$ when $n$ is even. Theorem 1.6 can be used to construct a double Gelfand model for $D_{n}$, namely, a character of $D_{n}$ in which every irreducible character appears exactly twice, see Theorem 7.3 below.
The rest of the paper is organized as follows. Section 2 contains necessary background and notation. In Section 3 we compute a generating function (Theorem 3.1) for root enumerators in $B_{n}$. Section 4 is devoted to higher Lie characters of type $B$, including an explicit definition (Definition 4.2) and a proof of a generating function (Theorem 4.3) for their values. In Section 5 we compare the two generating functions and prove Theorem 1.5, restated as Theorem 5.3. In Section 6 we consider the groups $D_{n}$ and prove Theorems 1.6, 1.2 and 1.8. The paper concludes with applications and open problems regarding Gelfand models (Subsection 7.1), index $2$ subgroups (Subsection 7.2), and related quasisymmetric functions (Subsection 7.3).
Acknowledgments
The authors thank Axel Hultman, Eric Marberg and Richard Stanley for useful discussions and references. They also thank the anonymous referee for his comments.
@@ 2.2. The Weyl groups of types $A$, $B$ and $D$
In this paper we use an explicit description of the conjugacy classes in $B_{n}$ (but do not need the corresponding description in $D_{n}$). A bipartition ${\underline{\lambda}}=(\lambda^{+},\lambda^{-})$ of a positive integer $n$, denoted ${\underline{\lambda}}\vdash n$, is a pair of partitions of total size $n$. Conjugacy classes in $B_{n}$ are in one-to-one correspondence with bipartitions of $n$. Each element $w\in B_{n}$, viewed as a permutation of the set [±$[\pm n]$, can be written as a product of disjoint cycles of total length $2n$. Moreover, if $c=(a_{1},a_{2},\ldots,a_{\ell})$ is such a cycle then so is $\bar{c}=(-{a}_{1},-{a}_{2},\ldots,-{a}_{\ell})$, and either $c$ and $\bar{c}$ are disjoint or else $c=\bar{c}$. In the former case, the product $c\bar{c}$ is said to be a positive cycle of $w$, of length $\ell$; otherwise $\ell$ is necessarily even and $c=\bar{c}$ is said to be a negative cycle of $w$, of length $\ell/2$. Then the pair of partitions $(\lambda^{+},\lambda^{-})$ for which the parts of $\lambda^{+}$ (respectively, $\lambda^{-}$) are the lengths of the positive (respectively, negative) cycles of $w$ is a bipartition of $n$, called the signed cycle type of $w$. Two elements of $B_{n}$ are conjugate if and only if they have the same signed cycle type. We denote by $C_{\underline{\lambda}}$ the conjugacy class of elements of $B_{n}$ of signed cycle type ${\underline{\lambda}}=(\lambda^{+},\lambda^{-})$.
@@ 2.3. Higher Lie characters of type $A$
$\psi_{S_{n}}^{\lambda}:=\omega^{\lambda}\uparrow_{Z_{\lambda}}^{S_{n}}.$
@@
$\rho^{S_{n}}_{k}=\sum_{\lambda\vdash_{k}n}\psi_{S_{n}}^{\lambda}.$
@@ 3. Generating function for root enumerators of type $B$
In this section we compute, for any integer $k$, an explicit generating function for the values of the $k$-th root enumerator $\rho_{k}^{B_{n}}$ on arbitrary elements of $B_{n}$. The result is expressed in terms of the $h$-th root enumerators $\rho_{h}^{{\mathbb{Z}}_{2}}$, corresponding to the cyclic group ${\mathbb{Z}}_{2}$ of order 2 (written multiplicatively as $\{+1,-1\}$) for various divisors $h$ of $k$.
@@ 3.1. Main result
$\sum_{n\geq 0}\sum_{{\underline{\mu}}\vdash n}\rho_{k}^{B_{n}}({\underline{\mu}})\,\frac{{\underline{t}}^{{\underline{c}}({\underline{\mu}})}}{|Z_{\underline{\mu}}|}=\exp\left(\sum_{j\geq 1}\sum_{\theta\in{\mathbb{Z}}_{2}}\sum_{\begin{subarray}{c}h|k\\ \gcd(h,j)=1\end{subarray}}\rho_{h}^{{\mathbb{Z}}_{2}}(\theta)\,\frac{t_{j,\theta}^{jk/h}}{2jk/h}\right).$
@@
$\sum_{n\geq 0}\sum_{\mu\vdash n}\rho_{k}^{S_{n}}(\mu)\,\frac{{\underline{t}}^{{\underline{c}}(\mu)}}{|Z_{\mu}|}=\exp\left(\sum_{j\geq 1}\sum_{\begin{subarray}{c}h|k\\ \gcd(h,j)=1\end{subarray}}\frac{t_{j}^{jk/h}}{jk/h}\right).$
@@
In order to compute the size of the conjugacy class of $y$, first write an arbitrary permutation of the numbers $1,\ldots,jb$. Interpret the first $j$ numbers as forming a cycle, and assign to each of the numbers an element of ${\mathbb{Z}}_{2}$ such that the ${\mathbb{Z}}_{2}$-class of the whole cycle is $\theta$. This can be done in $|{\mathbb{Z}}_{2}|^{j-1}$ ways. Treat similarly the other cycles, each having length $j$ and ${\mathbb{Z}}_{2}$-class $\theta$. Noting that any permutation of the $b$ cycles, as well as an arbitrary cyclic shift of each cycle, yields the same element of $B_{jb}$, it follows that the size of the conjugacy class of $y$ in $B_{jb}$ is
@@
so that $|R_{h,\theta}|=\rho^{{\mathbb{Z}}_{2}}_{h}(\theta)$.
@@
$\sum_{{\underline{\lambda}}\vdash n}\sum_{\begin{subarray}{c}x\in C_{\underline{\lambda}}\\ x^{k}=y\end{subarray}}1=\rho_{k}^{B_{n}}(y).$
Recall that $|R_{h,\theta}|=\rho_{h}^{{\mathbb{Z}}_{2}}(\theta)$ and $|Z_{\underline{\mu}}|=|B_{n}|/|C_{\underline{\mu}}|$. Denoting ${\underline{t}}^{{\underline{c}}({\underline{\mu}})}:=\prod_{j,\theta}t_{j,\theta}^{jb_{j,\theta}}$, Theorem 3.6 reduces to Theorem 3.1, presenting the (unrefined) generating function for the $k$-th root enumerator in $B_{n}$. ∎